Two projectiles are launched into the air. The first projectile has a y-component of its initial velocity of 10 m/s and an x-component of its initial velocity of 10 m/s. (This is equivalent to an initial speed of 14.2 m/s at a 45 degree angle.) The second projectile also has a y-component of its initial velocity of 10 m/s but an x-component of 5 m/s for its initial velocity. (This is equivalent to an initial speed of 11.2 m/s at a 63 degree angle.)

First, if both projectiles were launched at the same time on a horizontal surface, which one would land first, or would they land at the same time?

Second, which projectile would travel farther before landing, or would they travel the same distance?

Third, which projectile would reach the greatest maximum height off the ground, or would the maximum height be the same?
Answer #1

a)

from kinematic equation

Y=Yo+Voyt-(1/2)gt2

When projectile lands on ground Y=0

0=0+Voyt-(1/2)gt2

t=sqrt[2*Voy/g]

for first projectile

t1=sqrt(2*10/9.8) =1.43 s

for second projectile

t2=sqrt(2*10/9.8)=1.43 s

Therefore both projectile would land at the same time.

b)

Horizontal distance traveled by each projectile

D=Voxt

For each projectile

D1=10*1.43=14.3 m

D2=5*1.43=7.15 m

Therefore Projectile 1 would travel farther before landing.

C)

From Kinematic equation

Vfy2=Voy2-2gh

At maximum height h=Hmax ,Vfy=0

=>0=Voy2-2gHmax

Hmax =Voy2/2g

For each projectile

HMax,1=102/2*9.8 =5.1 m

HMax,2=102/2*9.8 =5.1 m

Therefore Both projectile would reach the same maximum height.